# DualSpec: Text-to-spatial-audio Generation via Dual-Spectrogram Guided Diffusion Model

### v1

@@ I Introduction
Spatial audio generation is essential for immersive extended reality (XR) environments, interactive entertainment systems, and dynamic media production. With the fast development of Artificial Intelligence Generated Content, there is a fancy question whether we could generate spatial audio from simply text that describes the sound events and spatial directions, known as text-to-spatial-audio (TTSA) generation. TTSA is a brand new research direction. It is rooted in the hot research area of text-to-audio (TTA) generation, which is a task of creating monaural audio from text descriptions. Its core challenge is how to guarantee the synthesis quality and azimuth accuracy of the spatial sound events simultaneously. To address this issue, TTA and conventional spatial audio generation techniques are involved. We summarize the two kinds of techniques as follows.
On the TTA side, the challenging issue of TTA is how to generate any kinds of sound events, covering from natural environments to human speech, flexibly with guaranteed high quality. Some recent development is as follows. AudioLM [1] utilizes the w2v-BERT model [2] to extract semantic tokens from audio inputs. AudioLDM [3] generates text-conditioned audio using a latent diffusion model (LDM) [4], where the diffusion process is guided by CLAP embeddings [5] and utilizes a variational autoencoder (VAE) [6] to generate latent embeddings of the Mel spectrograms of the audio. Tango [7] is built on AudioLDM [1]. It replaces CLAP with a fine-tuned large language model (LLM) FLAN-T5 [8]. Tango2 [9] further improved the performance of Tango by employing direct preference optimization [10] and alignment training.
@@ III DualSpec
In this section, we will introduce the proposed DualSpec framework, starting with its workflow pipeline, followed by a detailed explanation of the components it comprises, including the text encoder, diffusion model, VAE, and vocoder.
@@ III-B Text encoder
FLAN-T5-Small (80M) [8] is used as the text encoder ($E_{\text{text}}$), which generates text encodings $\tau\in\mathbb{R}^{L\times d_{\text{text}}}$, where $L$ is the token length and $d_{\text{text}}$ is the embedding dimension. Through pretraining on chain-of-thought and instructional datasets [29], FLAN-T5 gains the ability to effectively utilize in-context cues and mimic gradient-based optimization through its attention mechanisms. By treating each input as a separate task and leveraging its advanced pretraining, the model excels at extracting task-relevant information while reducing noise, ultimately enabling more accurate conversion from text to acoustic representations. We freeze the text encoder, which not only saves computational resources but also effectively prevents overfitting.
@@ III-D Classifier-free guidance
III-E Audio VAE and vocoder
Figure 2: The structural design of Dual-VAE demonstrates how to integrate STFT and Mel spectrogram features into a unified latent space representation.
@@ III-E Audio VAE and vocoder
Additionally, in order to explore the diverse combinations of the two acoustic features, we also trained the Dual-VAE model, which takes the concatenation of Mel spectrograms and STFT spectrograms as input and outputs their joint reconstructions, as illustrated in Fig. 2.
We use HiFi-GAN [34] as the vocoder to convert the Mel spectrogram into an audible waveform.
@@ V-C Data processing pipeline
First, we resampled each mono audio clip in Section V-A to a sample rate of 16kHz. and then we removed all audio clips that were shorter than 1s. In this way, we obtained a total of more than 90,000 mono audio clips. Next, we used HRIRS to convert each mono audio clip into corresponding dual-channel stereo audio files. We divided the horizontal plane into 12 directions, each separated by 30 degrees. Consequently, each mono audio clip was transformed into 12 stereo audio files, resulting in a total of approximately 1.2 million spatial stereo audio files.
For VAE training, we construct the training batch by randomly selecting one of the twelve spatial audio files associated with each audio event, cropping the audio length to 5 seconds. If the original audio length is less than 5 seconds, it is zero-padded to ensure the required length is met.
@@ VI-A Implementation details of VAEs
For the training of the reconstruction of the STFT spectrogram, we configured the VAE with a compression level of 8. We set the number of FFT points and the window length to 512, while the hop length was set to 256. For the reconstruction of the Mel spectrogram, we set the compression level to 4 and extracted 64 bands Mel spectrogram. When training the combination of two features, we followed the aforementioned settings and set the compression level to 8. We utilized the Adam optimizer [40] with a learning rate of $4.5\times 10^{-6}$ to train all VAE models. The training dataset, as described in Section V-C, consisted of 1.2 million samples. During each epoch, for every individual sound event, one out of the 12 available spatial audio was randomly selected for training. All models were trained for 800 epochs using four NVIDIA A100 40GB GPUs. The batch size was set to 64 for Mel and STFT features, and to 2 for the two-feature combination. To stabilize the training process, adversarial loss was not applied during the first 40 epochs.
We trained three VAEs with different acoustic features: Mel-VAE, STFT-VAE and Dual-VAE. As their names suggest, Mel-VAE and STFT-VAE were trained on Mel spectrograms and STFT spectrograms, respectively. The Dual-VAE, as described in Section III-E, was trained on both types of acoustic features simultaneously.
@@ VI-B Implementation details of diffusion models
Our diffusion model was built upon the UNet architecture from Stable Diffusion [4]. In this model, we configured it with 8 channels and a cross-attention dimension of 512. Pre-trained weights [9], which were trained on the AudioCaps mono audio dataset [41], were used. We used the AdamW optimizer [42] with a learning rate of $3\times 10^{-5}$ and a weight decay of $1\times 10^{-2}$. Additionally, a linear learning rate scheduler [43] was employed to adjust the learning rate over the course of training. We trained the diffusion model for 120 epochs, using four NVIDIA A100 40GB GPUs, with each GPU having a batch size of 2. Gradient accumulation was set to 8, resulting in an effective total batch size of 64. We set the number of inference steps to 200 and the classifer-free guidance to 3.
We have designed five different TTSA models, namely Mel-base, STFT-base, DualSpec, DualSpec-D, and DualSpec-P. Among them, Mel-base and STFT-base denoted diffusion models trained with Mel features and STFT features, respectively. They utilized Mel-VAE and STFT-VAE to compress acoustic features, respectively. DualSpec trained the diffusion model using a combination of latent representations encoded by the Mel-VAE and STFT-VAE, which is illustrated in Fig. 1. DualSpec-D utilized the latent representations encoded by Dual-VAE in Fig. 2 to train the diffusion model. DualSpec-P adopted an architecture similar to DualSpec, but utilized the VAE in [3] to encode Mel feature. Notably, this VAE can only process single-channel audio. Therefore, during the training process, we used VAE to encode Mel features on each channel of the spatial audio separately, and then concatenated the latent representations.
@@ VI-C1 Generation Quality Evaluation
We employ a series of metrics to comprehensively evaluate the generation quality of our VAEs and diffusion models. These include the Peak Signal-to-Noise Ratio (PSNR), which measures the noise level between the reconstructed and original signals, with higher values indicating better quality; the Structural Similarity Index (SSIM), assessing the similarity in structure between the reconstruction and the original content, where values closer to $1$ signify high-quality reconstruction; the Frechet Distance (FD), quantifying the discrepancy between generated results and real data, with lower values indicating more accurate reconstructions; the Inception Score (IS), evaluating the diversity and clarity of outputs from generative models, with higher scores reflecting superior generation quality; and the Kullback-Leibler Divergence (KL), measuring the difference between two probability distributions, where lower values indicate that the distribution of generated results closely matches that of the real data.
@@ VI-C2 Spatial Perception Evaluation
For spatial performance assessment, we employ standard sound source localization metrics from [35]: DOA Mean Absolute Error (MAE) and classification accuracy (ACC). While ACC represents a straightforward classification success rate, the circular nature of angular measurements on the azimuth plane necessitates special consideration for MAE calculations. Given a ground-truth angle $\theta$ and its corresponding prediction $\hat{\theta}$ for a sample, the MAE computation follows this angular distance formulation:
$\mathrm{MAE}(\circ)=\mathrm{min}(|\hat{\theta}-\theta|_{1},360-|\hat{\theta}-\theta|_{1})$ (17)
@@ VI-D Evaluation of VAEs using different features
TABLE I: Evaluation results for different VAE reconstruction performances.
@@ VI-E Comparison of different generation methods
      Baseline       AudioLDM [3]       21.45       2.470       7.194       -       -
      AudioLDM2 [44]       29.19       2.425       8.554       -       -
      TANGO [7]       27.40       2.919       9.572       -       -
      TANGO2 [9]       31.64       2.910       9.835       -       -
      Ours       Mel-base       15.47       2.058       10.691       32.28       24.54
      STFT-base       24.72       2.388       7.578       9.56       84.78
      DualSpec       23.87       2.376       7.590       13.16       79.02
      DualSpec-D       21.70       2.384       6.152       17.32       76.01
      DualSpec-P       20.49       2.322       10.571       10.79       83.86
@@ VII Conclusions
In conclusion, the proposed DualSpec demonstrates significant potential in TTSA generation, offering a novel approach that eliminates the need for monaural audio inputs. Under the guidance of text with azimuthal information, DualSpec achieves high directional consistency in generated spatial audio. Furthermore, the pipeline for the spatial audio dataset production also lays the foundation for TTSA tasks. Experimental results demonstrate the effectiveness of the proposed method in generating high-quality spatial audio with accurate spatial positioning.

### v2

@@ I Introduction
Spatial audio generation is essential for immersive extended reality (XR) environments, interactive entertainment systems, and dynamic media production. With the fast development of Artificial Intelligence Generated Content, there is a challenging question whether we could generate spatial audio from simply text that describes the sound events and spatial directions, known as text-to-spatial-audio (TTSA) generation. TTSA is a brand new research direction. It is rooted in the active research area of text-to-audio (TTA) generation, which is a task of creating monaural audio from text descriptions. Its core challenge is how to guarantee the synthesis quality and azimuth accuracy of the spatial sound events simultaneously. To address this issue, TTA and conventional spatial audio generation techniques are involved. We summarize the two kinds of techniques as follows.
On the TTA side, the challenging issue of TTA is how to generate any kinds of sound events, covering from natural environments to human speech, flexibly with guaranteed high quality. Some recent development is as follows. AudioLM [1] utilizes the w2v-BERT model [2] to extract semantic tokens from audio inputs. AudioLDM [3] generates text-conditioned audio using a latent diffusion model (LDM) [4], where the diffusion process is guided by CLAP embeddings [5] and utilizes a variational autoencoder (VAE) [6] to generate latent embeddings of the Mel spectrograms of the audio. [7] achieves significant improvements in semantic alignment and temporal consistency of generated audio by adopting a dual-text encoding architecture, a feed-forward Transformer-based diffusion denoiser and data augmentation strategies driven by large language models. Tango [8] is built on AudioLDM [3]. It replaces CLAP with a fine-tuned large language model (LLM) FLAN-T5 [9]. Tango2 [10] further improved the performance of Tango by employing direct preference optimization [11] and alignment training. Auffusion [12] leverages the advantages of text-to-image diffusion models in terms of generation capability and cross-modal alignment, significantly improving the quality of audio generation as well as the matching accuracy between text and audio under limited data and computational resources.
@@ III DualSpec
In this section, we will introduce the proposed DualSpec framework, starting with its workflow pipeline, followed by a detailed explanation of the components it comprises, including the text encoder, diffusion model and VAE.
@@ III-B Text encoder
FLAN-T5-Large [9] is used as the text encoder ($E_{\text{text}}$), which generates text encodings $\tau\in\mathbb{R}^{L\times d_{\text{text}}}$, where $L$ is the token length and $d_{\text{text}}$ is the embedding dimension. Through pretraining on chain-of-thought and instructional datasets [31], FLAN-T5 gains the ability to effectively utilize in-context cues and mimic gradient-based optimization through its attention mechanisms. By treating each input as a separate task and leveraging its advanced pretraining, the model excels at extracting task-relevant information while reducing noise, ultimately enabling more accurate conversion from text to acoustic representations. We freeze the text encoder, which not only saves computational resources but also effectively prevents overfitting.
@@ III-D Classifier-free guidance
III-E Audio VAE
Figure 2: The structural design of Dual-VAE demonstrates how to integrate STFT and Mel spectrogram features into a unified latent space representation. The dashed line indicates that this process occurs only during the inference phase.
@@ III-E Audio VAE
Additionally, in order to explore the diverse combinations of the two acoustic features, we also trained the Dual-VAE model, which takes the concatenation of Mel spectrograms and STFT spectrograms as input and outputs their joint reconstructions, as illustrated in Fig. $2$ The Inverse-transform in Fig. 2 is implemented using least squares optimization to minimize the Euclidean distance between the Mel spectrogram and the product of the estimated magnitude spectrogram and the filter banks. .
@@ V-C Data processing pipeline
First, we resampled each mono audio clip in Section V-A to a sample rate of 16kHz, and then we removed all audio clips that were shorter than 1s. In this way, we obtained a total of more than 90000 mono audio clips. Next, we used HRIRS to convert each mono audio clip into corresponding dual-channel stereo audio files. We divided the horizontal plane into 12 directions, each separated by 30 degrees. Consequently, each mono audio clip was transformed into 12 stereo audio files, resulting in a total of approximately 1.2 million spatial stereo audio files.
For VAE training, we construct each training batch using the spatial audio files mentioned above, cropping the audio length to 5 seconds. If the original audio length is less than 5 seconds, it is zero-padded to ensure the required length is met.
@@ VI-A Implementation details of VAEs
For the training of the reconstruction of the STFT spectrogram, we configured the VAE with a compression level of 8. We set the number of FFT points and the window length to 512, while the hop length was set to 256. For the reconstruction of the Mel spectrogram, we set the compression level to 4 and extracted 64 bands Mel spectrogram. When training the combination of two features, we followed the aforementioned settings and set the compression level to 8. We utilized the Adam optimizer [41] with a learning rate of $4.5\times 10^{-6}$ to train all VAE models. The training dataset, as described in Section V-C, consisted of 1.2 million samples.
We trained three VAEs with different acoustic features: Mel-VAE, STFT-VAE and Dual-VAE. As their names suggest, Mel-VAE and STFT-VAE were trained on Mel spectrograms and STFT spectrograms, respectively. The Dual-VAE, as described in Section III-E, was trained on both types of acoustic features simultaneously. All VAE models were trained on 16 NVIDIA H100 GPUs, with approximately 0.6 million training steps for Mel-VAE, 1.2 million for STFT-VAE, and 2.4 million for Dual-VAE. The batch sizes per GPU were set to 16, 8, and 4 for Mel-VAE, STFT-VAE, and Dual-VAE, respectively. To stabilize the training process, adversarial loss was not applied during the first 60 thousand steps.
@@ VI-B Implementation details of diffusion models
Our diffusion model was built upon the UNet architecture from Stable Diffusion [4]. In this model, we configured it with 8 channels and a cross-attention dimension of 1024. Pre-trained weights [10], which were trained on the AudioCaps mono audio dataset [42], were used. We used the AdamW optimizer [43] with a learning rate of $3\times 10^{-5}$ and a weight decay of $1\times 10^{-2}$. Additionally, a linear learning rate scheduler [44] was employed to adjust the learning rate over the course of training. We trained the diffusion model for 200 epochs, using 4 NVIDIA H100 GPUs, with each GPU having a batch size of 2. Gradient accumulation was set to 8, resulting in an effective total batch size of 64. We set the number of inference steps to 200 and the classifer-free guidance to 3. We employed the pre-trained FLAN-T5-Large [9] model as the text encoder and keep its parameters frozen during the training process.
We have designed five different TTSA models, namely Mel-base, STFT-base, DualSpec, DualSpec-D, and DualSpec-P. Among them, Mel-base and STFT-base denoted diffusion models trained with Mel features and STFT features, respectively. They utilized Mel-VAE and STFT-VAE to compress acoustic features, respectively. DualSpec trained the diffusion model using a combination of latent representations encoded by the Mel-VAE and STFT-VAE, which is illustrated in Fig. 1. DualSpec-D utilized the latent representations encoded by Dual-VAE in Fig. 2 to train the diffusion model. DualSpec-P adopted an architecture similar to DualSpec, but utilized the VAE in [3] to encode Mel feature. Notably, this VAE can only process single-channel audio. Therefore, during the training process, we used VAE to encode Mel features on each channel of the spatial audio separately, and then concatenated the latent representations. For Mel-base, DualSpec, and DualSpec-P, we used the HiFi-GAN vocoder [45] with pre-trained weights from [3] to convert the Mel spectrogram into an audible waveform.
@@ VI-C1 Generation Quality Evaluation
We employ the following metrics to evaluate the generation quality of our VAEs and diffusion models:
  • Peak Signal-to-Noise Ratio (PSNR): It reflects the extent of signal distortion by quantifying the difference between the original and test signal. Its mathematical definition is as follows:
    $\text{PSNR}=10\cdot\log_{10}\left(\frac{\text{MAX}^{2}}{\text{MSE}}\right),$ (17)
    where MAX represents the maximum possible value of the audio signal, and MSE denotes the mean squared error between the original audio and the test audio.
  • Structural Similarity Index (SSIM): It evaluates the degree of similarity by analyzing the correlations between signals in terms of luminance, contrast, and structure. Specifically, it is implemented by comparing the mean, variance, and covariance of the signals, with its definition as follows:
    $\text{SSIM}=\frac{(2\mu_{x}\mu_{y}+c_{1})(2\sigma_{xy}+c_{2})}{(\mu_{x}^{2}+\mu_{y}^{2}+c_{1})(\sigma_{x}^{2}+\sigma_{y}^{2}+c_{2})},$ (18)
    where $u_{x}$ and $u_{y}$ represent the mean values of the test signal and the reference signal, respectively. $\sigma_{x}^{2}$ and $\sigma_{y}^{2}$ are their corresponding variances, and $\sigma_{xy}$ represents the covariance between the two signals. $c_{1}$ and $c_{2}$ are constant parameters related to the dynamic range of the audio.
  • Frechet Distance (FD): It is a metric commonly used to assess the quality of generated audio by quantifying the discrepancy between the feature distributions of real and generated audio. Specifically, audio signals are first encoded into feature vectors using the PANNs audio feature extraction model [46], after which the FD score is computed according to the following formula:
    $\text{FD}=\|\bar{\mu}_{y}-\bar{\mu}_{x}\|^{2}+\text{Tr}\left(\Sigma_{y}+\Sigma_{x}-2\cdot\sqrt{\Sigma_{y}\Sigma_{x}}\right),$ (19)
    where $\bar{\mu}_{x}$ and $\bar{\mu}_{y}$ denote the mean vectors of the generated and real audio features across all dimensions, respectively, while $\Sigma_{x}$ and $\Sigma_{y}$ represent the corresponding covariance matrices.
  • Kullback-Leibler Divergence (KL): It is used to quantify the discrepancy between the probability distributions of real and generated audio. Before computing the divergence, audio features are extracted using the PANNs model. A smaller KL divergence indicates that the distribution of the generated audio is closer to that of the real audio.
  • Inception Score (IS): It assesses both the quality and diversity of generated samples, and its calculation is defined as follows:
    $\text{IS}=\exp\left(\mathbb{E}_{x}\left[\text{KL}(p(y|x)\parallel p(y))\right]\right),$ (20)
    where $p(y|x)$ represents the classification probability distribution for a single audio sample, $p(y)$ denotes the marginal distribution over all samples, and $\text{KL}(A\parallel B)$ indicates the KL score between distributions A and B. The pre-trained classification model is PANNs.
@@ VI-C2 Spatial Perception Evaluation
For spatial performance assessment, we employ standard sound source localization metrics from [36], which are detailed as follows:
  • Classification Accuracy (ACC): It measures the proportion of generated audio samples whose spatial directions are consistent with those of the corresponding real audio. The training data includes audio from 12 different DOAs, with audio from each DOA categorized into a separate class. We use the sound source localization model from Section IV to obtain the DOA labels for the generated audio. The ACC is formulated as follows:
    $\text{ACC}=\frac{N_{\text{correct}}}{N_{\text{total}}},$ (21)
    where $N_{\text{correct}}$ denotes the number of generated audio clips whose DOA matches that of the corresponding real audio clips, and $N_{\text{total}}$ denotes the total number of generated audio clips.
  • Mean Absolute Error (MAE): It is a commonly used metric for evaluating the performance of direction estimation models. Due to the periodic nature of directional angle values, MAE is adopted as an appropriate metric to quantify the difference between the DOAs of the generated and real audio, thereby improving the reliability of the evaluation. Given a ground-truth angle $\theta$ and its corresponding prediction $\hat{\theta}$ for a sample, the MAE computation follows this angular distance formulation:
    $\mathrm{MAE}(\circ)=\mathrm{min}(|\hat{\theta}-\theta|_{1},360-|\hat{\theta}-\theta|_{1}).$ (22)
@@ VI-D Evaluation of VAEs using different features
TABLE I: Evaluation results for different VAE reconstruction performances. $\downarrow$ indicates that lower values are better, and vice versa for $\uparrow$.
@@ VI-E Comparison of different generation methods
Baseline AudioLDM [3] 809 452 0.36 21.45 2.470 7.194 - -
AudioLDM2 [47] 1034 537 0.71 29.19 2.425 8.554 - -
TANGO [8] 1317 646 0.53 27.40 2.919 9.572 - -
TANGO2 [10] 1317 646 0.52 31.64 2.910 9.835 - -
Ours Mel-base 1485 679 0.65 15.47 2.058 10.691 32.28 24.54
STFT-base 1262 848 0.97 24.72 2.388 7.578 9.56 84.78
DualSpec 1541 1145 1.21 23.87 2.376 7.590 13.16 79.02
DualSpec-D 1262 887 1.06 21.70 2.384 6.152 17.32 76.01
DualSpec-P 1373 1417 1.82 20.49 2.322 10.571 10.79 83.86
@@
Table II also presents a comparison of the computational load across models, including parameter sizes, FLOPs for one second of audio, and inference time for the same duration. Compared to the baseline models, DualSpec and its variants show a relatively significant increase in both FLOPs and inference time. The main reasons are as follows. First, due to the use of two features rather than relying solely on the Mel spectrogram as in the baseline method, DualSpec incurs higher computational costs. Second, spatial audio contains more data than ordinary mono audio, which also increases computational costs to some extent.
@@ VII Conclusions
In conclusion, the proposed DualSpec demonstrates significant potential in TTSA generation, offering a novel approach that eliminates the need for monaural audio inputs. In addition, DualSpec effectively mitigates the limitations inherent in relying on a single feature, as the Mel spectrogram exhibits low azimuth accuracy and the STFT spectrogram leads to degraded audio quality. We also introduce spatial perception metrics to evaluate the azimuth accuracy of the generated spatial audio. Furthermore, the pipeline for the spatial audio dataset production also lays the foundation for TTSA tasks. Experimental results demonstrate that the proposed method can generate spatial audio that balances both audio quality and directional consistency.